# HyperFusion: A Hypernetwork Approach to Multimodal Integration of Tabular and Medical Imaging Data for Predictive Modeling

### v1

@@ 4.1.5 Results and discussion
A comparison of the hypernetwork framework to the training of two separate networks each with approximately half the data (either males or females) clearly demonstrates that the HyperFusion enables a much more efficient data utilization. Integrating information from subjects of both sexes, the hypernetwork enhanced the prediction accuracy as shown in Figure 5B. Although not within the scope of this paper, it is interesting to note that female data age prediction is much more accurate than male age prediction (see left-most and middle bars in Figure 5as A and B). This could be derived from the higher variability of age related brain appearance in males.
4.2 Alzheimer’s disease classification
To further assess the efficiency of our network in fusing imaging and tabular data, we conducted a set of experiments on the AD classification task conditioned by different number and compositions of tabular data attributes. We compared our results to those obtained using either of the data modalities as well as to other existing approaches.
@@ 4.2.1 The data
We used four phases of the ADNI database (ADNI1, ADNI2, ADNI GO, and ADNI3) and sorted out datasets of 2120 subjects. Since part of the ADNI data is longitudinal we excluded part of the data to ensure each subject has a single, yet complete, visit’s dataset, i.e., a single MRI scan and the (temporally) corresponding EHR. Statistics of the ADNI data used are presented in Table 1. The concrete imaging and EHR datasets we used along with the corresponding ADNI’s subject identification numbers are available at the Git repository enclosed to this work.
Table 1: Statistics on the CN, MCI and AD groups in our data. Information includes the number of samples ”N” per label, the mean and standard deviation of age, and the male/female ratio within each diagnostic group
@@ 4.2.4 Data partitioning
The dataset is partitioned into five non-overlapping subsets of equal size, ensuring that each subset maintains a similar joint distribution of age, sex, and diagnosis, in accordance with the methodology outlined by (Wen et al., 2020). Among these five subsets, four are allocated for cross-validation, while the fifth is set aside for testing, as depicted in Appendix B. The partitioning involves a random shuffling of the data using a predefined random seed, referred to as the split seed. This systematic approach to data splitting enhances the validity and generality of the evaluation process, showcasing the robustness of our model.
@@ 4.2.5 Ablation study and comparisons
The imaging-tabular fusion models we compared include various methods. The first one, dubbed ‘concatenation’, involved employing our pre-activation ResNet backbone (presented in Figure 4B) yet, replacing the HyperRes block with a regular pre-activation Res block. The tabular attributes were concatenated with the second-to-last fully connected layer, a technique documented in prior works like (Esmaeilzadeh et al., 2018).
Additionally, we compared our method to the results reported for two contemporary imaging-tabular fusion techniques: the DAFT (Wolf et al., 2022) and the late-fusion approach proposed by (Prabhu et al., 2022). To ensure a fair comparison and highlight the unique (isolated) contribution of our hypernetwork, we constructed and trained two additional networks. Specifically, we utilized our primary network backbone and training regimen, including our proposed regularized weighted categorical loss function, maintaining the methodologies of the FiLM (Perez et al., 2018) (FiLM-like implementation) and the DAFT (Wolf et al., 2022) (DAFT-like implementation). We note our implementation of the FiLM was adapted from (Wolf et al., 2022), since originally the FiLM was designed for text and natural image fusion.
Figure 6: Bar plot presentation of the AD classification results for six competing models and ours using seven different metrics (Section 4.2.6). As previously noted, we included the TP (True Positive) rate for each class as a metric, namely CN (Cognitively Normal), MCI (Mild Cognitive Impairment), and AD (Alzheimer’s Disease). The metrics are represented along the X-axis, with each metric featuring a set of bars corresponding to all competing models. The Y-axis presents the scores achieved by each model for each metric. The bars present the mean and standard deviation over the test data in all random splits and versions. The compared methods including ‘$*{}^{*}$’ and ‘$*{}^{*}$’ are described in Section 4.2.5.
Table 2: AD classification results for the proposed method and other competing models. The table presents a comparison of the AD classification task using the HyperFusion method (first row) with respect to baseline methods using either tabular data (second row) or imaging data (third row) alone. The results of a straightforward imaging-tabular concatenation strategy, as well as our improved implementation of the DAFT-like and the FiLM-like models are presented in the 4-th, 5-th and 6-th rows respectively. The last rows present reported results of the late-fusion method of (Prabhu et al., 2022) and the DAFT (Wolf et al., 2022). The comparison is based on seven metrics detailed in Section 4.2.6. As previously noted, we included the TP (True Positive) rate for each class as a metric, namely CN (Cognitively Normal), MCI (Mild Cognitive Impairment), and AD (Alzheimer’s Disease). For our method and our implementation of other models we report the mean and standard deviation over the test data in all random splits and versions. In addition, for each metric we calculated the p-value from a Mann-Whitney U test comparing our method to the others to assess the statistical significance of the differences.
@@
(Prabhu et al., 2022)
$*{}^{*}$our adapted implementation as described in Section 4.2.5
@@ 4.2.6 Training and evaluation
We utilized the Adam optimizer for training, coupled with a regularized weighted cross-entropy loss (Equations 1, 3), as outlined in Sections 3.2 and 3.5.2. The embedding MLP, discussed in detail in Section 3.5.2 and depicted in Figure 4C, underwent pre-training exclusively with the training tabular data to expedite the convergence of the entire network compound.
As mentioned in Section 3.5.2, each subject’s imaging data includes two 3D subimages of the left and right hippocampus. During training, we augmented the dataset by randomly selecting either subimage in each feed-forward iteration. For improved predictions during validation and testing, we processed each subject’s data twice, once for each brain side, and averaged the soft decisions. Furthermore, we achieved enhanced test results by employing an ensemble model, as elaborated in Section 3.4.
To guarantee the robustness of our findings, we adopted an extensive cross-validation strategy encompassing three unique split seeds, each undergoing three rounds of random initialization (versions). In every cross-validation iteration, four models were trained and validated on distinct data folds. For effective result utilization, the four models from a specific version and split seed were aggregated to assess their collective performance on a shared, unseen test set. A visualization for better understanding the training, validation, and testing processes is provided in a figure at Appendix B.
We employed various metrics to assess the performance of our framework and compare it with other models. These metrics include balanced accuracy (BA) and precision (PRC), defined as follows:
$\mbox{BA}=\frac{1}{3}{\sum_{c}}\frac{TP_{c}}{TP_{c}+FN_{c}}~{}~{}~{}~{}~{}~{}~{}\mbox{PRC}=\frac{1}{3}{\sum_{c}}\frac{TP_{c}}{TP_{c}+FP_{c}}$ (7)
@@ 4.2.7 Results and discussion
In this section, we conduct a comprehensive comparison between all the models discussed in Section 4.2.5 and our hypernetwork model, assessing both global and class-specific metrics. The results are summarized in Table 2 and Figure 6. It is noteworthy that all multi-modal methods consistently outperformed unimodal models in terms of global performance metrics and in nearly all class-specific metrics.
Table 2 presents the reported scores for DAFT (Wolf et al., 2022) and a late fusion method (Prabhu et al., 2022) for certain metrics, along with the results obtained for the methods implemented by us as described in Section 4.2.5. For a fair evaluation of the comparison, we note the following differences.
The late fusion method of (Prabhu et al., 2022) used datasets of 3,256 subjects (35.4% CN, 51.1% MCI, 13.5% AD) from the same ADNI phases that we utilized. The tabular data contained an extensive set of 20 attributes, including patient demographics, medical history, vital signs, neurophysiological test results, cognitive function, and other relevant diagnostic information. Some of these attributes were deliberately excluded from our study, as noted in 4.2.3. Notably, the late fusion method obtained higher PRC scores at the expense of weaker BA results due to the inherent trade-off between these two metrics using imbalanced datasets. By applying class weight adjustments to the loss function during training, we were able to achieve better PRC and BA compared to the reported late fusion method results, despite using fewer subjects and a smaller set of tabular attributes (results are shown in the Appendix D). The reported DAFT method (Wolf et al., 2022) used datasets of 1,341 subjects (40.3% CN, 40.1% MCI, 19.6% AD) from the same ADNI source and a similar set of tabular attributes. To address the difference in the data size, we presented the scores obtained using our implementation of a DAFT-like method (Section 4.2.5) as well.
@@ 5 Discussion and Conclusions
In summary, this research not only identifies and tackles a crucial deep-learning challenge but also provides a robust solution through hypernetworks. The potential implications of this work extend from improving diagnostics to shaping the broader landscape of personalized medical care. As DNNs in the medical domain continue their rapid evolution, methods like ours, bridging the gap between diverse data modalities, are poised to play a significant role.
@@ Acknowledgments
Appendix A Brain Age Prediction Data
[…]
@@ Appendix A Brain Age Prediction Data
UK Biobank (Sudlow et al., 2015) 17,088 63.1 (±7.5) 8,118:8,970
Consortium for Reliability and Reproducibility, CoRR (Zuo et al., 2014) 1,350 26.1 (±15.9) 661:689
Alzheimer’s Disease Neuroimaging Initiative, ADNI (Jack Jr et al., 2008) 659 72.9 (±6.0) 432:227
Brain Genomics Superstruct Project, GSP (Buckner et al., 2014) 1,099 21.5 (±2.9) 469:630
Functional Connectomes Project, FCP (Biswal et al., 2010) 922 28.2 (±13.9) 393:529
Autism Brain Imaging Data Exchange, ABIDE (Di Martino et al., 2014) 506 16.9 (±7.7) 411:95
Parkinson’s Progression Markers Initiative, PPMI (Marek et al., 2011) 172 60.3 (±11.0) 105:67
International Consortium for Brain Mapping, ICBM (Mazziotta et al., 1995) 592 30.9 (±12.5) 319:273
Australian Imaging, Biomarkers and Lifestyle, AIBL (Ellis et al., 2009) 604 73.0 (±6.6) 268:335
Southwest University LongitudinalImaging Multimodal, SLIM (Liu et al., 2017) 571 20.1 (±1.3) 251:320
Information extraction from Images, IXI (Heckemann et al., 2003) 562 48.2 (±16.5) 251:311
Open Access Series of Imaging Studies, OASIS (Marcus et al., 2007, 2010) 401 51.7 (±24.9) 144:257
Consortium for Neuropsychiatric Phenomics, CNP (Poldrack et al., 2016) 124 31.5 (±8.8) 65:59
Center for Biomedical Research Excellence, COBRE (Mayer et al., 2013) 74 35.8 (±11.6) 51:23
Child and Adolescent NeuroDevelopment Initiative, CANDI (Frazier et al., 2008) 29 10.0 (±2.9) 17:12
Brainomics (Pinel et al., 2012) 68 25.4 (±7.1) 31:37
Cambridge Centre for Ageing and Neuroscience, CamCAN (Shafto et al., 2014) 652 54.3 (±18.6) 322:330
Child Mind Institute - Healthy Brain Network, HBN (Alexander et al., 2017) 304 10.0 (±3.7) 164:140
Nathan Kline Institute-Rockland Sample, NKI (Nooner et al., 2012) 914 37.1 (±21.7) 359:555
@@
Appendix B AD data splitting, model training and evaluation
Figure 7 visualize the data partitioning and well as the evaluation method for the cross validation experiments of the AD classification.
Figure 7: Data Splitting and model training and evaluation illustration for AD Classification Task: each version is a different random initialization of a model. The entire process is performed using three random seeds with three versions for each, i.e., nine cross-validation experiments all together.
Appendix C AD Classification Confusion Matrices
Figure 8 presents the confusion matrices for our hyperfusion and the compared methods for the AD classification task, providing more insight on the relative classification abilities for each subgroup.
@@ Appendix C AD Classification Confusion Matrices
Appendix D Ablation Study: Class Weights for the AD Classification Task
As mentioned in the main manuscript, we trained the networks with weighted categorical cross entropy loss. The weights of the classes influence the relative scores between the different classes. The late fusion method (Prabhu et al., 2022) obtained higher PRC scores at the expense of weaker BAC results due to the inherent trade-off between these two metrics using imbalanced datasets. We note that for other combinations of class weights our HyperFusion method achieved higher scores in all the reported metrics. Table 4 presents the results obtained by our method, trained with different class weight configurations as well as the reported results of the late fusion method to support our claim.
Table 4: AD classification results for the proposed HyperFusion and the late-fusion method. The first row presents the scores obtained our model that trained using different class weight configurations in the weighted cross entropy loss function. The second row presents reported results of the late-fusion method of (Prabhu et al., 2022). The comparison is based on three metrics detailed in main manuscript.

### v2

@@ 4.2.6 Results and discussion
The comparisons presented in Figure 4 clearly show that the proposed hyperfusion model outperforms the image-only network for all male-female train-test configurations, as well as the standard concatenation framework. Although beyond the scope of this paper, it is interesting to note that the prediction of female age is much more accurate than the prediction of male age (see the leftmost and middle bars in Figure 4A and B). We assume that this might be due to the greater variability of brain age in males.
4.3 Multi-class AD classification
To further assess the efficiency of our network in fusing imaging and tabular data, we performed a set of experiments on the multi-class AD classification task conditioned by different number and compositions of tabular data attributes. We compared our results to those obtained using either of the data modalities as well as to other existing approaches.
@@ 4.3.1 The data
We used four phases of the ADNI databases (ADNI1, ADNI2, ADNI GO, and ADNI3) - selecting a single and complete dataset of each individual (obtained during the first visit). Each of the selected datasets includes a single T-1 MRI scan and the (temporally) corresponding EHR. Statistics of the ADNI data used are presented in Table 2. The information includes the number of samples ”N” per label, the mean and standard deviation of age, and the male/female ratio within each diagnostic group. The data used along with the corresponding ADNI’s subject identification numbers are available at the Git repository enclosed with this work.
Table 2: Statistics on the CN, MCI and AD sub-populations in our data.
@@ 4.3.4 Data partitioning
The dataset is partitioned into five non-overlapping subsets of equal size, ensuring that each subset maintains a similar joint distribution of age, sex, and diagnosis, in accordance with the methodology outlined by [67]. Among these five subsets, four are allocated for cross-validation, while the fifth is set aside for testing, as depicted in C.1. The partitioning involves a random shuffling of the data using a predefined random seed, referred to as the split seed. This systematic approach to data splitting enhances the validity and generality of the evaluation process, showcasing the robustness of our model.
@@ 4.3.5 Ablation study and comparisons
The imaging-tabular fusion models we compared include various methods. The first one, dubbed ‘concatenation’, involved employing our pre-activation ResNet backbone (presented in Figure 3B) yet, replacing the HyperRes block with a regular pre-activation Res block. The tabular attributes were concatenated with the second-to-last fully connected layer, a technique documented in prior works like [16].
Additionally, we compared our method to the results reported for two contemporary imaging-tabular fusion techniques: the DAFT [68] and the late-fusion approach proposed by [53]. To ensure a fair comparison and highlight the unique (isolated) contribution of our hypernetwork, we constructed and trained two additional networks. Specifically, we utilized our primary network backbone and training regimen, including our proposed regularized weighted categorical loss function, maintaining the methodologies of FiLM [49] (FiLM-like implementation) and DAFT [68] (DAFT-like implementation). We note that our implementation of the FiLM was adapted from [68], since originally the FiLM was designed for text and natural image fusion.
Figure 5: Bar plot presentation of the AD classification results for six competing models and ours using seven different metrics (Section 4.3.6). The compared methods including ‘DAFT-like’ and ‘FiLM-like’ are described in Section 4.3.5.
Table 3: AD classification results for the proposed method and other competing models using seven different metrics.
@@
[53]
our adapted implementation as described in Section 4.3.5
@@ 4.3.6 Training and evaluation
We use the Adam optimizer for training, coupled with a regularized WCE loss (Equations 3, 5), as outlined in Sections 3.2 and 3.5.2. The WCE loss accounts for the imbalance between the CN, MCI and AD classes in the training data. However, using standard categorical cross entropy loss while over-sampling the less frequent classes can handle the class imbalance problem as well. The embedding MLP, discussed in detail in Section 3.5.2 and depicted in Figure 3C, underwent pre-training exclusively with the training tabular data to expedite the convergence of the entire network compound.
As mentioned in Section 3.5.2, each subject’s imaging data includes two 3D subimages of the left and right hippocampus. During training, we augment the dataset by randomly selecting either subimage in each feed-forward iteration. For better predictions during validation and testing, we processed each subject’s data twice, once for each brain side, and averaged the soft decisions. Furthermore, we achieved enhanced test results employing an ensemble model, as elaborated in Section 3.4.
To guarantee the robustness of our findings, we adopted an extensive cross-validation strategy encompassing three unique split seeds, each undergoing three rounds of random initialization (versions). In every cross-validation iteration, four models were trained and validated on distinct data folds. For effective result utilization, the four models from a specific version and split seed were aggregated to assess their collective performance on a shared, unseen test set. A visualization for better understanding the training, validation, and testing processes is provided in Figure 7 in C.1.
We used various metrics to assess the performance of our framework and compare it with other models. These metrics include balanced accuracy (BA) and precision (PRC), defined as follows:
${\color[rgb]{0,0,0}\mbox{BA}=\frac{1}{3}{\sum_{c}}\frac{TP_{c}}{TP_{c}+FN_{c}}}$ (9)
${\color[rgb]{0,0,0}\mbox{PRC}=\frac{1}{3}{\sum_{c}}\frac{TP_{c}}{TP_{c}+FP_{c}}}$ (10)
@@ 4.3.7 Results and discussion
In this section, we present a comprehensive comparison between all the models discussed in Section 4.3.5 and our hypernetwork model, evaluating both global and class-specific metrics. The results are summarized in Table 3 and Figure 5. In both the bar plot and the table we present the mean and standard deviation over the test data in all random splits and versions. The compared methods that include ‘DAFT-like’ and ‘FiLM-like’ are described in Section 4.3.5.
It is noteworthy that all multi-modal methods consistently outperformed unimodal models in terms of global performance metrics and in nearly all class-specific metrics.
Table 3 also presents the scores reported for DAFT [68] and a late fusion method [53] for certain metrics, along with the results obtained for the methods we implemented as described in Section 4.3.5. For a fair evaluation of the comparison, we note the following differences.
The late fusion method of [53] used datasets of 3,256 subjects (35.4% CN, 51.1% MCI, 13.5% AD) from the same ADNI phases that we utilized. The tabular data contained an extensive set of 20 attributes, including patient demographics, medical history, vital signs, neurophysiological test results, cognitive function, and other relevant diagnostic information. Some of these attributes were deliberately excluded from our study, as noted in 4.3.3. Notably, the late fusion method obtained higher PRC scores at the expense of weaker BA results due to the inherent trade-off between these two metrics using imbalanced datasets. By applying class weight adjustments to the loss function during training, we were able to achieve better PRC and BA compared to the reported late fusion method results, despite using fewer subjects and a smaller set of tabular attributes (results are shown in C.4). The reported DAFT method [68] used datasets of 1,341 subjects (40.3% CN, 40.1% MCI, 19.6% AD) from the same ADNI source and a similar set of tabular attributes. To address the difference in data size, we presented the scores obtained using our implementation of a DAFT-like method (Section 4.3.5) as well.
@@ 5 Discussion and Conclusions
In summary, this research not only identifies and tackles a crucial deep-learning challenge but also provides a robust solution through hypernetworks. The potential implications of this work extend from improving diagnostics to shaping the broader landscape of personalized medical care. As DNNs in the medical domain continue their rapid evolution, methods like ours, which bridge the gap between diverse data modalities, are poised to play a significant role.
@@ Acknowledgments
Appendix A Brain Age Prediction
A.1 Datasets and Preprocessing
Table 4 presents the datasets used for the brain age prediction task. The preprocessing pipeline of the T1-weighted MRI scans is identical to the one proposed in [Levakov et al., 2020], using Nipype [Gorgolewski et al., 2011]. Initially, RobustFov Jenkinson et al. [2012] was employed to remove the neck and shoulders from each scan. Subsequently, Robex Iglesias et al. [2011] was utilized for brain extraction, followed by expanding the brain mask to encompass the CSF surrounding the brain. The correction for intensity non-uniformity was carried out using N4BiasFieldCorrection Tustison et al. [2010]. Further, intensity normalization was implemented using fuzzy c-means and WM-based mean normalization Reinhold et al. [2018]. Ultimately, the images were resampled to a resolution of 1.75 mm3 and cropped to a 90x120x99 voxels box centered around the brain mask’s center of mass.
@@ A.1 Datasets and Preprocessing
UK Biobank [Sudlow et al., 2015] 17,088 63.1 (±7.5) 8,118:8,970
Consortium for Reliability and Reproducibility, CoRR [Zuo et al., 2014] 1,350 26.1 (±15.9) 661:689
Alzheimer’s Disease Neuroimaging Initiative, ADNI [Jack Jr et al., 2008] 659 72.9 (±6.0) 432:227
Brain Genomics Superstruct Project, GSP [Buckner et al., 2014] 1,099 21.5 (±2.9) 469:630
Functional Connectomes Project, FCP [Biswal et al., 2010] 922 28.2 (±13.9) 393:529
Autism Brain Imaging Data Exchange, ABIDE [Di Martino et al., 2014] 506 16.9 (±7.7) 411:95
Parkinson’s Progression Markers Initiative, PPMI [Marek et al., 2011] 172 60.3 (±11.0) 105:67
International Consortium for Brain Mapping, ICBM [Mazziotta et al., 1995] 592 30.9 (±12.5) 319:273
Australian Imaging, Biomarkers and Lifestyle, AIBL [Ellis et al., 2009] 604 73.0 (±6.6) 268:335
Southwest University LongitudinalImaging Multimodal, SLIM [Liu et al., 2017] 571 20.1 (±1.3) 251:320
Information extraction from Images, IXI [Heckemann et al., 2003] 562 48.2 (±16.5) 251:311
Open Access Series of Imaging Studies, OASIS [Marcus et al., 2007, 2010] 401 51.7 (±24.9) 144:257
Consortium for Neuropsychiatric Phenomics, CNP [Poldrack et al., 2016] 124 31.5 (±8.8) 65:59
Center for Biomedical Research Excellence, COBRE [Mayer et al., 2013] 74 35.8 (±11.6) 51:23
Child and Adolescent NeuroDevelopment Initiative, CANDI [Frazier et al., 2008] 29 10.0 (±2.9) 17:12
Brainomics [Pinel et al., 2012] 68 25.4 (±7.1) 31:37
Cambridge Centre for Ageing and Neuroscience, CamCAN [Shafto et al., 2014] 652 54.3 (±18.6) 322:330
Child Mind Institute - Healthy Brain Network, HBN [Alexander et al., 2017] 304 10.0 (±3.7) 164:140
Nathan Kline Institute-Rockland Sample, NKI [Nooner et al., 2012] 914 37.1 (±21.7) 359:555
@@
A.2 Brain Age Prediction Model Architecture
Table 5 presents the architectural details of the proposed HyperFusion model for brain age prediction presented in Section 3.5.1 of the main paper.
Table 5: Architecture of the brain age prediction model
Brain Age Prediction Model Architecture
Index Name Type Params
0 conv1_a Conv3d 448
1 conv1_b Conv3d 6.9 K
2 batchnorm1 BatchNorm3d 32
3 conv2_a Conv3d 13.9 K
4 conv2_b Conv3d 27.7 K
5 batchnorm2 BatchNorm3d 64
6 conv3_a Conv3d 55.4 K
7 conv3_b Conv3d 110 K
8 batchnorm3 BatchNorm3d 128
9 dropout1 Dropout3d 0
10 linear1 HyperLinear 1.3 M
11 linear2 HyperLinear 1.1 K
12 linear3 HyperLinear 4.2 K
13 final_layer HyperLinear 133
Trainable params 1.5 M
Total params 1.5 M
Total estimated model params size (MB) 5.929
Appendix B Binary AD classification
We adapted our hypernetwork framework which performs multi-class Alzheimer Disease (AD) classification to address binary classification, thus further demonstrating the strength of our method. The binary classification task is considered less challenging since structural brain differences between AD patients and cognitively normal (CN) subjects are apparent. Nevertheless, there is still a significant gain in integrating both imaging and tabular data using the proposed hypernetwork framework on our sorted ADNI dataset, as is demonstrated in Figure 6. To further demonstrate the advantage of the proposed hypernetwork framework we fused the soft classification results of both the tabular-only and the imaging-only and present the resulting late-fusion predictions (orange bars) in the figure. While the late-fusion results are better than those obtained by the tabular-only for all metrics and imaging-only for most metrics they are inferior with respect to the hypernetwork performances.
Figure 6: Binary classification experiments comparing imaging/tabular data single modality and Hyperfusion.
Appendix C Multiclass AD classification
C.1 Data splitting, model training and evaluation
Figure 7 depicts the partition of the ADNI data as well as the evaluation method for the cross validation experiments of the AD classification.
Figure 7: A schematic illustration of the data splitting and model training and evaluation for multiclass AD classification: each round is based on a different random initialization of a model. The entire process is performed using three random seeds with three versions for each, i.e., nine cross-validation experiments all together.
C.2 AD Classification Model Architecture
Table 6 presents the architectural details of the proposed HyperFusion model for AD classification presented in Section 3.5.2 of the main paper.
Table 6: Architecture of the AD classification model
AD Classification Model Parameters
Index Name Type Params
0 conv_bn_relu Sequential 480
1 max_pool3d_1 MaxPool3d 0
2 block1 PreactivResBlock_bn 42.2 K
3 block2 PreactivResBlock_bn 168 K
4 block3 PreactivResBlock_bn 672 K
5 block4 HyperPreactivResBlock 3.0 M
6 adaptive_avg_pool3d AdaptiveAvgPool3d 0
7 linear_drop1 Dropout 0
8 fc1 LinearLayer 16.4 K
9 linear_drop2 Dropout 0
10 fc2 LinearLayer 195
11 relu ReLU 0
Trainable params 3.9 M
Total params 3.9 M
Total estimated model params size (MB) 15.415
C.3 AD Classification Confusion Matrices
Figure 8 presents the confusion matrices obtained by the proposed the hyperfusion model for the AD in comparison to other methods. The confusion matrices provide statistics on the the accurate and inaccurate classification rates of each class.
@@ C.3 AD Classification Confusion Matrices
C.4 Class Weights for the AD Classification Task
As mentioned in the main manuscript, we trained the hyperfusion framewotk for multi-class AD classification with weighted categorical cross entropy (WCE) loss, where the weights of the classes (CN, MCI, AD) are approximately inverse-proportional to their frequency. To assess our choice, we performed an ablation study comparing the proposed approach with a prevalent alternative to address data imbalance, an oversampling of the less frequent classes using standard CE loss. The results are presented in Table 7 showing the advantage of using the WCE loss.
Table 7: Hyperfusion performances using WCE/CE loss with/without upsampling
model BA PRC f1 macro AUC macro TP-CN TP-MCI TP-AD
WCE loss .673 .624 .630 .822 .759 .495 .764
CE w upsample .643 .597 .601 .806 .687 .488 .755
CE w/o upsample .628 .595 .599 .782 .668 .521 .695
Table 8: AD classification results for the proposed HyperFusion and the late-fusion method. The first row presents the scores obtained our model using different class weight configurations in the weighted cross entropy loss function. The second row presents reported results of the late-fusion method of [Prabhu et al., 2022]. The comparison is based on three metrics detailed in main manuscript.